# Solving dynamic portfolio selection problems via score-based diffusion models

### v1

@@ 3 Stability of mean-variance portfolio selection problems
Suppose the stock prices $(S^{1},S^{2},\cdots,S_{T})$ satisfies some model $\mathbb{P}$. Suppose the risk-free interest rate is 0. For a investment strategy $\vartheta$ (in some admissible or constraint set), the value associated with it is $(\vartheta\cdot S)_{t}:=\sum_{l=1}^{t-1}\vartheta_{l}^{\mathbf{t}}(S_{l+1}-S_{l})$. We want to maximize the mean-variance objective:
@@ 3.1 Technical preparations
Denote $\Pi^{S}_{t}:=\sqrt{\mathbb{E}_{\mathbb{P}}[\triangle S^{t}(\triangle S^{t})^{\mathbf{t}}|\mathcal{F}_{t}]},t=1,2,\cdots,T-1$. Then for any $\vartheta,\bar{\vartheta}\in L^{2}(S)$, $(\vartheta\cdot S)=(\bar{\vartheta}\cdot S)$ if and only if $\Pi^{S}\vartheta=\Pi^{S}\bar{\vartheta}$, $t=1,2,\cdots,T-1$.
@@
From definition we know $(\vartheta\cdot S)=(\bar{\vartheta}\cdot S)$ if and only if for any $t=1,2,\cdots,T-1$, $\vartheta^{\mathbf{t}}\triangle S^{t}=\bar{\vartheta}^{\mathbf{t}}\triangle S^{t}$, which in turn is equivalent to
@@
We only need to prove the closedness. Suppose $(\vartheta^{n}\cdot S)_{T}\to(\vartheta\cdot S)_{T}$ in $L^{2}(\mathbb{P})$ for some $\vartheta^{n}\in\Theta_{K,\mathbb{P}}$, and $\vartheta\in L^{2}(S)$. Consider the semimartingale decomposition of $S$, $S^{t}=s_{0}+A_{t}+M_{t}$, where $\triangle M_{t}:=S^{t+1}-\mathbb{E}_{\mathbb{P}}[S^{t+1}|\mathcal{F}_{t}]$, $\triangle A_{t}:=\mathbb{E}_{\mathbb{P}}[\triangle S^{t}|\mathcal{F}_{t}]$, $t=1,2,\cdots,T-1$. For any $t=1,2,\cdots,T-1$, we have
@@
Starting from $t=T-1$, because $(\vartheta^{n}\cdot S)_{T}\to(\vartheta\cdot S)_{T}$ in $L^{2}(\mathbb{P})$, we know $\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\triangle M_{T-1}\big{|}^{2}\Big{]}\to 0$. From (3.6), it is not hard to derive
$\mathbb{E}_{\mathbb{P}}[|\Pi^{S}_{T-1}(\vartheta^{n}_{T-1}-\vartheta_{T-1})|^{2}]=$ $\mathbb{E}_{\mathbb{P}}\Big{[}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\mathbb{E}_{\mathbb{P}}[\triangle S_{T-1}\triangle S_{T-1}^{\mathbf{t}}|\mathcal{F}_{T-1}](\vartheta^{n}_{T-1}-\vartheta_{T-1})\Big{]}$ (3.17)
@@
$=\frac{\delta}{1-\delta}\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}(\vartheta^{n}_{t}-\vartheta_{t})^{\mathbf{t}}\triangle M_{t}\big{|}^{2}\Big{]}$ (3.19)
@@
Thus $\Pi^{S}_{T-1}\vartheta^{n}_{T-1}\to\Pi^{S}_{T-1}\vartheta_{T-1}$ in $L^{2}(\mathbb{P})$. Equivalently, we have $(\vartheta^{n}_{T-1})^{\mathbf{t}}\triangle S_{T-1}\to\vartheta_{T-1}^{\mathbf{t}}\triangle S_{T-1}$ in $L^{2}(\mathbb{P})$. By using (3.15) and repeating the arguments, we actually conclude that $\Pi^{S}_{t}\vartheta^{n}_{t}\to\Pi^{S}_{t}\vartheta_{t}$ in $L^{2}(\mathbb{P})$ for any $t=1,2,\cdots,T-1$, hence up to a subsequence, the convergence is almost surely. Because $K$ is closed convex, we can define the projection operator $\Pi_{K}:\mathbb{R}^{d}\to K$. Clearly $\Pi^{S}_{t}\Pi_{K}(\mathbb{R}^{d})$ is a closed subset of $\mathbb{R}^{d}$, $\mathbb{P}$-almost surely. Therefore, there exists a $\vartheta^{\prime}$ such that $\Pi^{S}_{t}\Pi_{K}\vartheta^{\prime}_{t}=\Pi^{S}_{t}\vartheta_{t}$. By Lemma 3.3, $(\vartheta\cdot S)=(\Pi_{K}\vartheta^{\prime}\cdot S)$, thus $(\vartheta\cdot S)_{T}\in G(\Theta_{K,\mathbb{P}})$. ∎
@@
There exists an optimizer in $\Theta(K)$ to the quadratic hedging problem (3.3), and the optimizer is unique in value.
@@ 3.2 Dynamic programming
and for the initial problem $V(1,0,s^{1})$, an optimizer is given recursively by $\hat{\vartheta}_{t}=\hat{\theta}(t,(\hat{\vartheta}\cdot S)_{t},S_{1:t})$, where
@@
(3.23) is clear. Because the optimizer of (3.21) exists, we pick one such optimizer $\hat{\vartheta}_{l}\in\Theta^{t,s^{1:t}}_{K,\mathbb{P}}$, $l=t,t+1,\cdots,T-1$. By the definition of $\mathcal{F}^{t}_{l}$, $\hat{\vartheta}_{l}$ can be written as deterministic functions $\theta(l,S^{t+1},\cdots,S^{l})$. For any $s^{t+1}$, the random variable $\tilde{\vartheta}^{s^{t+1}}_{l}$ given by $\omega\mapsto\theta(l,s^{t+1},S^{t+2},\cdots,S^{l})$ is $\mathcal{F}^{t+1}_{l}$ measurable. Thus, $\{\tilde{\vartheta}_{l}^{s^{t+1}}\}_{l=t+1}^{T-1}\in\Theta^{t+1,s_{1:t+1}}_{K,\mathbb{P}}$. Consequently,
@@
$\geq\inf_{\theta_{t}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\inf_{\theta_{t+1}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t},S^{t+1}}}\Big{[}V\Big{(}t+2,w+\theta_{t}^{\mathbf{t}}\triangle S^{t}+\theta_{t+1}^{\mathbf{t}}\triangle S^{t+1},(s^{1:t},S^{t+1},S_{t+2})\Big{)}\Big{]}$ (3.30)
$\geq\inf_{\theta_{t},\theta_{t+1}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}t+2,w+\theta_{t}^{\mathbf{t}}\triangle S^{t}+\theta_{t+1}^{\mathbf{t}}\triangle S^{t+1},(s^{1:t},S^{t+1},S_{t+2})\Big{)}\Big{]}$ (3.31)
[…]
@@ 3.3 Stability
Denote by $V(\cdot;c,\mathbb{Q})$ the function obtained by DPP (3.23)-(3.24) with $\mathbb{P}$ replaced by $\mathbb{Q}$. The next theorem aims to bound $|V-\tilde{V}|$ by $\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})$.
@@
For simplicity, we use $V$ and $\tilde{V}$ to denote $V(\cdot;c,\mathbb{P})$ and $V(\cdot;c,\mathbb{Q})$, respectively. We prove (3.38) by induction. $t=T$ is clear from the terminal condition. Suppose we have (3.38) for $t+1$. Because $\pi\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$, $\pi_{s^{1:t},\tilde{s}^{1:t}}\in\Pi(\mathbb{P}_{s^{1:t}},\mathbb{Q}_{\tilde{s}^{1:t}})$. Therefore, by considering a near-optimal $\vartheta$ in (3.24), it is not hard to show that
@@
$-\tilde{V}(t+1,\tilde{w}+\vartheta^{\mathbf{t}}(\tilde{S}^{t+1}-\tilde{s}_{t}),(\tilde{s}^{1:t},\tilde{S}^{t+1}))\big{|}$ (3.41)
Note that in this proof, we adopt the notation $S\sim\mathbb{P}$, $\tilde{S}\sim\mathbb{Q}$, and the lower case letters represent realized (sampled) paths, which are used in regular conditional kernels. From Assumption 4, we know $\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+1}|^{2}$ is bounded, uniformly in $s^{1:t}$. Moreover, $\vartheta\in K$ is uniformly bounded by assumption. We thus derive from (3.41) and the induction hypothesis that
@@
$\times$ $\Big{(}\mathbb{E}_{\pi_{S_{1:t+1},\tilde{S}_{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}+\mathbb{E}_{\pi_{S_{1:t+1},\tilde{S}_{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Bigg{]}$ (3.45)
=:absent:=:= : $J_{1}+J_{2}+J_{3}.$ (3.46)
@@
×$\times\Big{(}\mathbb{E}_{\pi_{S_{1:t+1},\tilde{S}_{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}$ (3.54)
@@
$+C\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Bigg{[}\Big{|}S^{t+1}-\tilde{S}^{t+1}\Big{|}\cdot\Big{(}\mathbb{E}_{\pi_{S_{1:t+1},\tilde{S}_{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}\Bigg{]}$ (3.56)
@@
Therefore, the maximum in (3.61) is attained in a bounded interval $[0,\tilde{L}]$, and we conclude from (3.39) that
@@ 4.1 Implementation of Algorithm 1
Here, for $t=1,2,\cdots,T$, $\{s^{1:t}_{(m)}\}_{m=1}^{M}$ is a batch of data, and $\{h^{1:t}_{(m)}\}_{m=1}^{M}$ are the corresponding hidden variables obtained by feeding the data to the RNN encoder $R_{\theta}$; $\{\tau_{(k)}\}_{k=1}^{K}$ are randomly sampled from $[0,\mathcal{T}]$; $\{\mathbf{z}_{n}\}_{n=1}^{N}$ are sampled from pure noise $\mathcal{N}(0,I)$. Although only $s_{\theta}$ explicitly appears in (4.1), $R_{\theta}$ is implicitly trained by backward propagation. Algorithm 2 provides the pseudocode for the training procedure.
Algorithm 2 Training the score network and the feature extraction module
@@
Outputs: Trained score network $s_{\theta}$ and feature extraction module $R_{\theta}$.
As for the sampling procedure, we consider an implementable version of Algorithm 1, which has the following two refinements: we utilize the hidden variables obtained from the feature extraction module to resolve high dimensionality; we also use the conventional predictor-corrector sampler to reduce the discretization error from the SDE solver. The sampling procedure can be summarized in Algorithm 3. Note that Algorithm 3 is the implementation version of Algorithm 1.
@@
Inputs: Trained score network $s_{\theta}$ and feature extraction module $R_{\theta}$; number time discretization (predictor) steps $\mathrm{N_{pre}}$; number of corrector steps $\mathrm{N_{cor}}$; predictor steps $\{\tau_{k}\}_{k=1}^{\mathrm{N_{pre}}}$; corrector step sizes $\{\mathbf{e}_{k}\}_{k=1}^{\mathrm{N_{pre}}}$; initial feature $h^{0}$
1:  $h\leftarrow h^{0}$.

### v3

@@ 3 Stability of mean-variance portfolio selection problems
Suppose the stock prices $(S^{1},S^{2},\cdots,S^{T})$ satisfies some model $\mathbb{P}$. Suppose the risk-free interest rate is 0. For a investment strategy $\vartheta$ (in some admissible or constraint set), the value associated with it is $(\vartheta\cdot S)_{t}:=\sum_{l=1}^{t-1}\vartheta_{l}^{\mathbf{t}}(S^{l+1}-S^{l})$. We want to maximize the mean-variance objective:
@@ 3.1 Technical preparations
Denote $\Pi^{S}_{t}:=\sqrt{\mathbb{E}_{\mathbb{P}}[\triangle S^{t}(\triangle S^{t})^{\mathbf{t}}|\mathcal{F}_{t}]},t=1,2,\cdots,T-1$. Then for any $\vartheta,\bar{\vartheta}\in L^{2}(S)$, $(\vartheta\cdot S)_{t}=(\bar{\vartheta}\cdot S)_{t}$ for any $t=1,2,\cdots,T-1$ if and only if $\Pi^{S}_{t}\vartheta=\Pi^{S}_{t}\bar{\vartheta}$ for any $t=1,2,\cdots,T-1$.
@@
From definition we know $(\vartheta\cdot S)_{t}=(\bar{\vartheta}\cdot S)_{t}$ for any $t=1,2,\cdots,T-1$ if and only if $\vartheta^{\mathbf{t}}\triangle S^{t}=\bar{\vartheta}^{\mathbf{t}}\triangle S^{t}$, which in turn is equivalent to
@@
We only need to prove the closedness. Suppose $(\vartheta^{n}\cdot S)_{T}\to(\vartheta\cdot S)_{T}$ in $L^{2}(\mathbb{P})$ for some $\vartheta^{n}\in\Theta_{K,\mathbb{P}}$, and $\vartheta\in L^{2}(S)$. Consider the semimartingale decomposition of $S$, $S^{t}=s^{1}+A_{t}+M_{t}$, where $\triangle M_{t}:=S^{t+1}-\mathbb{E}_{\mathbb{P}}[S^{t+1}|\mathcal{F}_{t}]$, $\triangle A_{t}:=\mathbb{E}_{\mathbb{P}}[\triangle S^{t}|\mathcal{F}_{t}]$, $t=1,2,\cdots,T-1$. For any $t=1,2,\cdots,T-1$, we have
@@
Starting from t=T1t=T-1, because $(\vartheta^{n}\cdot S)_{T}\to(\vartheta\cdot S)_{T}$ in $L^{2}(\mathbb{P})$, we know $\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\triangle M_{T-1}\big{|}^{2}\Big{]}\to 0$. From (3.6), it is not hard to derive
$\mathbb{E}_{\mathbb{P}}[|\Pi^{S}_{T-1}(\vartheta^{n}_{T-1}-\vartheta_{T-1})|^{2}]=$ $\mathbb{E}_{\mathbb{P}}\Big{[}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\mathbb{E}_{\mathbb{P}}[\triangle S^{T-1}(\triangle S^{T-1})^{\mathbf{t}}|\mathcal{F}_{T-1}](\vartheta^{n}_{T-1}-\vartheta_{T-1})\Big{]}$ (3.17)
@@
$=\frac{\delta}{1-\delta}\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\triangle M_{T-1}\big{|}^{2}\Big{]}$ (3.19)
@@
Thus $\Pi^{S}_{T-1}\vartheta^{n}_{T-1}\to\Pi^{S}_{T-1}\vartheta_{T-1}$ in $L^{2}(\mathbb{P})$. Equivalently, we have $(\vartheta^{n}_{T-1})^{\mathbf{t}}\triangle S^{T-1}\to\vartheta_{T-1}^{\mathbf{t}}\triangle S^{T-1}$ in $L^{2}(\mathbb{P})$. By using (3.15) and repeating the arguments, we actually conclude that $\Pi^{S}_{t}\vartheta^{n}_{t}\to\Pi^{S}_{t}\vartheta_{t}$ in $L^{2}(\mathbb{P})$ for any $t=1,2,\cdots,T-1$, hence up to a subsequence, the convergence is almost surely. Because $K$ is closed convex, we can define the projection operator $\Pi_{K}:\mathbb{R}^{d}\to K$. Clearly $\Pi^{S}_{t}\Pi_{K}(\mathbb{R}^{d})$ is a closed subset of $\mathbb{R}^{d}$, $\mathbb{P}$-almost surely. Therefore, there exists a $\vartheta^{\prime}$ such that $\Pi^{S}_{t}\Pi_{K}\vartheta^{\prime}_{t}=\Pi^{S}_{t}\vartheta_{t}$. By Lemma 3.3, $(\vartheta\cdot S)=(\Pi_{K}\vartheta^{\prime}\cdot S)$, thus $(\vartheta\cdot S)_{T}\in G(\Theta_{K,\mathbb{P}})$. ∎
@@
There exists an optimizer in $\Theta(K)$ to the quadratic hedging problem (3.3), and the optimizer is unique in value, i.e., $(\vartheta\cdot S)_{T}=(\tilde{\vartheta}\cdot S)_{T}$, $\mathbb{P}$-a.s., for any two optimizers $\vartheta$ and $\tilde{\vartheta}$.
@@ 3.2 Dynamic programming
and for the initial problem $V(1,0,s^{1})$, an optimizer is given recursively by $\hat{\vartheta}_{t}=\hat{\theta}(t,(\hat{\vartheta}\cdot S)_{t},S^{1:t})$, where
@@
(3.23) is clear. Because the optimizer of (3.21) exists, we pick one such optimizer $\hat{\vartheta}_{l}\in\Theta^{t,s^{1:t}}_{K,\mathbb{P}}$, $l=t,t+1,\cdots,T-1$. By the definition of $\mathcal{F}^{t}_{l}$, $\hat{\vartheta}_{l}$ can be written as deterministic functions $\theta(l,S^{t+1},\cdots,S^{l})$. For any $s^{t+1}$, the random variable $\tilde{\vartheta}^{s^{t+1}}_{l}$ given by $\omega\mapsto\theta(l,s^{t+1},S^{t+2},\cdots,S^{l})$ is $\mathcal{F}^{t+1}_{l}$ measurable. Thus, $\{\tilde{\vartheta}_{l}^{s^{t+1}}\}_{l=t+1}^{T-1}\in\Theta^{t+1,s^{1:t+1}}_{K,\mathbb{P}}$. Consequently,
@@
$\geq\inf_{\theta_{t}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\inf_{\theta_{t+1}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t},S^{t+1}}}\Big{[}V\Big{(}t+2,w+\theta_{t}^{\mathbf{t}}\triangle S^{t}+\theta_{t+1}^{\mathbf{t}}\triangle S^{t+1},(s^{1:t},S^{t+1},S^{t+2})\Big{)}\Big{]}$ (3.30)
$\geq\inf_{\theta_{t},\theta_{t+1}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}t+2,w+\theta_{t}^{\mathbf{t}}\triangle S^{t}+\theta_{t+1}^{\mathbf{t}}\triangle S^{t+1},(s^{1:t},S^{t+1},S^{t+2})\Big{)}\Big{]}$ (3.31)
$\geq\cdots$ (3.32)
@@ 3.3 Stability
Denote by $V(\cdot;c,\mathbb{Q})$ the function obtained by DPP (3.23)-(3.24) with $\mathbb{P}$ replaced by $\mathbb{Q}$, and we also write $V=V(\cdot;c,\mathbb{P})$ to emphasize the dependence on $c$ and $\mathbb{P}$. The next theorem aims to bound |VV~||V-$\tilde${V}| by $\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})$.
@@
For simplicity, we use $V$ and $\tilde{V}$ to denote $V(\cdot;c,\mathbb{P})$ and $V(\cdot;c,\mathbb{Q})$, respectively. We prove (3.38) by induction. $t=T$ is clear from the terminal condition. Suppose we have (3.38) for $t+1$. Because $\pi\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$, we have $\pi_{s^{1:t},\tilde{s}^{1:t}}\in\Pi(\mathbb{P}_{s^{1:t}},\mathbb{Q}_{\tilde{s}^{1:t}})$ (c.f. Backhoff et al. (2017), Proposition 5.1). Therefore, by considering a near-optimal $\vartheta$ in (3.24), it is not hard to show that
@@
$-\tilde{V}(t+1,\tilde{w}+\vartheta^{\mathbf{t}}(\tilde{S}^{t+1}-\tilde{s}^{t}),(\tilde{s}^{1:t},\tilde{S}^{t+1}))\big{|}$ (3.41)
Note that in this proof, we adopt the notation $\sim\mathbb{P}, S~\tilde{S}\sim\mathbb{Q}$, and the lower case letters represent realized (sampled) paths, which are used in regular conditional kernels. From Assumption 4, we know $\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+1}|^{2}$ is bounded, uniformly in $s^{1:t}$. Moreover, $\vartheta\in K$ is uniformly bounded by assumption. We thus derive from (3.41) and the induction hypothesis that
@@
$\times$ $\Big{(}\mathbb{E}_{\pi_{S^{1:t+1},\tilde{S}^{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}+\mathbb{E}_{\pi_{S^{1:t+1},\tilde{S}^{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Bigg{]}$ (3.45)
$=:$ $J_{1}+J_{2}+J_{3}.$ (3.46)
@@
$\times\Big{(}\mathbb{E}_{\pi_{S^{1:t+1},\tilde{S}^{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}$ (3.54)
@@
$+C\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Bigg{[}\Big{|}S^{t+1}-\tilde{S}^{t+1}\Big{|}\cdot\Big{(}\mathbb{E}_{\pi_{S^{1:t+1},\tilde{S}^{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}\Bigg{]}$ (3.56)
@@
Therefore, the maximum in (3.61) is attained in a bounded interval $0,\tilde{L}$], and we conclude from (3.39) that
@@
Corollary 3.9 ensures that, measured in adapted Wasserstein metric, MV value functions are stable in the deviation of models. Furthermore, the next corollary justifies that solving the problem with $\mathbb{Q}$ gives a approximated optimal solution to the original problem under $\mathbb{P}$.
Corollary 3.10.
Suppose $\vartheta^{*}_{\varepsilon}$ is a suboptimal solution to $v^{*}(\mathbb{Q})$ such that
$v(\mathbb{Q};\vartheta^{*}_{\varepsilon})\geq v^{*}(\mathbb{Q})-\varepsilon.$ (3.69)
Then, we have
$v(\mathbb{P};\vartheta^{*}_{\varepsilon})\geq v^{*}(\mathbb{P})-\varepsilon-C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}).$ (3.70)
Proof.
In the proof of Theorem 3.8 and Corollary 3.10, we can fixed a $\vartheta^{*}_{\varepsilon}\in\Theta_{K,\mathbb{Q}}$, and because $|\vartheta^{*}_{\varepsilon}|\leq K$, we have
$|v(\mathbb{P};\vartheta^{*}_{\varepsilon})-v(\mathbb{Q};\vartheta^{*}_{\varepsilon})|\leq C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}).$
Therefore, by triangle inequality
$|v(\mathbb{P};\vartheta^{*}_{\varepsilon})-v^{*}(\mathbb{P})|\leq$ $|v(\mathbb{P};\vartheta^{*}_{\varepsilon})-v(\mathbb{Q};\vartheta^{*}_{\varepsilon})|+|v(\mathbb{Q};\vartheta^{*}_{\varepsilon})-v^{*}(\mathbb{Q})|+|v^{*}(\mathbb{Q})-v^{*}(\mathbb{P})|$ (3.71)
$\leq$ $\varepsilon+2C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}),$ (3.72)
which gives (3.70) because $v(\mathbb{P};\vartheta^{*}_{\varepsilon})\leq v^{*}(\mathbb{P})$. ∎
@@ 4.1 Implementation of Algorithm 1
Here, $\{s^{1:T}_{(m)}\}_{m=1}^{M}$ is a batch of data, and $\{h^{1:T}_{(m)}\}_{m=1}^{M}$ are the corresponding hidden variables obtained by feeding the data to the RNN encoder $R_{\theta}$; $\{\tau_{(m)}\}_{m=1}^{M}$ are randomly sampled from $0,\mathcal{T}]$; $\{\mathbf{z}_{(m)}\}_{m=1}^{M}$ are sampled from pure noise $\mathcal{N}(0,I)$. Although only $s_{\theta}$ explicitly appears in (4.1), $R_{\theta}$ is implicitly trained by backward propagation. Algorithm 2 provides the pseudocode for the training procedure.
Algorithm 2 Training the score network and the RNN encoder
@@
Outputs: Trained score network $s_{\theta}$ and RNN encoder $R_{\theta}$.
As for the sampling procedure, we consider an implementable version of Algorithm 1, which has the following two refinements: we utilize the hidden variables obtained from the RNN encoder to resolve high dimensionality; we also use the conventional predictor-corrector sampler to reduce the discretization error from the SDE solver. The sampling procedure can be summarized in Algorithm 3.
@@
Inputs: Trained score network $s_{\theta}$ and RNN encoder $R_{\theta}$; number time discretization (predictor) steps $\mathrm{N_{pre}}$; number of corrector steps $\mathrm{N_{cor}}$; predictor steps $\{\tau_{k}\}_{k=1}^{\mathrm{N_{pre}}}$; corrector step sizes $\{\mathbf{e}_{k}\}_{k=1}^{\mathrm{N_{pre}}}$; initial feature $h^{0}$
[…]